# and thermofield double states as quantum many-body scars

### v3

@@ Abstract
We establish a duality between open many-body systems governed by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation and satisfying the detailed balance condition on the one side, and closed bilayer systems with a self-adjoint Hamiltonian on the other side. Under this duality, the identity operator on the open system side maps to the thermofield double state which turns out to be a quantum many-body scar of the dual Hamiltonian $\mathcal{H}$. A remarkable feature of this thermofield scar is a tunable entanglement entropy controlled by the reservoir temperature on the open system side. Further, we identify broad classes of many-body open systems with nontrivial explicit eigen operators $Q$ of the Lindbladian superoperator. The expectation values of the corresponding observables exhibit a simple exponential decay, $\langle Q\rangle_{t}=e^{-\Gamma t}\langle Q\rangle_{0}$, irrespectively of the initial state. Under the above duality, these eigen operators give rise to additional (towers of) scars. Finally, we point out that more general superoperators (not necessarily of the GKSL form) can be mapped to self-adjoint Hamiltonians of bilayer systems harbouring scars, and provide an example thereof.
@@
Remarkably, the scar obtained here turns out to be the thermofield double state, thus we adopt the term “thermofield scar”. The thermofield double state is a highly entangled state that first emerged in thermal quantum field theory Umezawa et al. (1982) and further found applications in molecular dynamics Borrelli and Gelin (2021), black hole physics Maldacena (2003); Maldacena and Susskind (2013); Hartman and Maldacena (2013) and quantum simulations Wu and Hsieh (2019); Zhu et al. (2020); Francis et al. (2021). Importantly, the bipartite entanglement entropy of the thermofield scar is thermal, with the temperature coinciding with the reservoir temperature on the open system side of the duality. In a special case of infinite temperature, the entanglement entropy is maximal and the thermofield scar reduces to the recently discovered Einstein-Podolsky-Rosen (EPR) scar Langlett et al. (2022); Wildeboer et al. (2022).
@@
The rest of this paper is organized as follows. We start from introducing the duality. Then we show that the duality implies the existence of the thermofield scar. After that we illustrate our general findings in a particular family of spin-$1/2$ models. Further, we demonstrate that in broad classes of dissipative models one can construct explicit eigen operators of the adjoint Lindbladian, and that these operators map into additional scars in dual bilayer closed systems. Then we briefly discuss dualities between bilayer Hamiltonians and superoperators of non-GKSL form. We conclude by discussing our results in relation to other developments and outlining further prospects of our approach.
@@
Here $O_{t}$ is a Heisenberg operator of some observable, ${\mathcal{L}}$ is a Lindbladian superoperator (Lindbladian), $H$ is the Hamiltonian which is responsible for the coherent part of evolution, and $\cal D$ is the dissipation superoperator (dissipator). The operator $O_{t}$ acts in a Hilbert space $\mathscr{H}$. The expectation value $\langle O\rangle_{t}$ of the observable $O$ evolves in time according to $\langle O\rangle_{t}=\operatorname{\mbox{tr}}\rho_{0}\,O_{t}$, where $\rho_{0}$ is the initial state of the open system.
It is well-known that an open system can be formally mapped to a closed system with doubled degrees freedom, i.e. with the Hilbert space $\mathscr{H}\otimes\mathscr{H}$. Under this map, any operator $O$ on the open system side maps to a vector $|O\rangle\rangle$ in $\mathscr{H}\otimes\mathscr{H}$, the Lindbladian superoperator $i{\mathcal{L}}^{\dagger}$ maps to the “Hamiltonian” operator $\cal H$ acting in $\mathscr{H}\otimes\mathscr{H}$, and the GKSL equation (1) maps to the “Schrödinger equation” $i\partial_{t}|O_{t}\rangle\rangle={\cal H}|O_{t}\rangle\rangle$ Baranger (1958); Havel (2003); Medvedyeva et al. (2016). This map is merely a homomorphism between two linear spaces, the space of operators on the open system side and the space of wave vectors on the closed system side. Importantly, under this map the operator $\cal H$ is never self-adjoint (except a trivial case of no dissipation, ${\cal D}=0$). While such non-Hermitian Hamiltonians can be useful Roccati et al. (2022), in general they do not correspond to any physical system.
@@
The latter condition naturally emerges when deriving the GKSL equation for a system interacting with a thermal reservoir in weak coupling or low density limits (Alicki and Lendi, 2007, Sec. 1.3.4). Under this condition, the dissipator has the form
@@
where $\beta$ is a real parameter having the meaning of the inverse reservoir temperature in microscopic derivations of the GKSL equation (Breuer et al., 2002, Sec. 3.3.2), $\gamma>0$ is the dissipative coupling constant and (rescaled) Lindblad operators $L_{j}$ are annihilation-like operators with respect to some $H_{0}$ that commutes with $H$,
$[\beta H_{0},L_{j}]=-\beta L_{j},\quad[H_{0},H]=0.$ (3)
@@
$+\mathbb{1}\otimes\left(H^{\mathsf{T}}-\frac{g}{4}\sum_{j}\left(e^{\frac{\beta}{2}}L_{j}^{\mathsf{T}}L_{j}^{*}+e^{-\frac{\beta}{2}}L_{j}^{*}L_{j}^{\mathsf{T}}\right)\right)$
$+\frac{g}{2}\sum_{j}\biggl{(}L_{j}^{\dagger}\otimes L_{j}^{\mathsf{T}}+L_{j}\otimes L_{j}^{*}\biggr{)}.$ (6)
@@
The duality (5), (6) is the first main result of the present paper. An infinite-temperature version of this duality was outlined in our previous paper Teretenkov and Lychkovskiy (2024).
@@
This is the second main result of the paper.
The state (8) is the thermofield double state Umezawa et al. (1982). In general, the double layer Hamiltonian (6) is nonintegrable and zero energy is in the middle of its spectrum, in which case the eigenstate (8) is classified as a scar Serbyn et al. (2021); Moudgalya et al. (2022). A particular case of the thermofield scar (8) with $\beta=0$ has been recently discovered in refs. Langlett et al. (2022); Wildeboer et al. (2022). [^1] See also ref. Cottrell et al. (2019) where a local bilayer Hamiltonian is constructed whose ground state is approximately the thermofield double state. It has been tagged as the EPR scar in ref. Wildeboer et al. (2022), for reasons that will be clarified in the example below.
If one traces out one layer of the thermofield scar (8), one obtains a reduced state of the second layer which turns out to be a Gibbs state with respect to $H_{0}$ (Nastase, 2023, Section 32.5), (sup, , Eq. (S8))
@@
where normalization is omitted. As a consequence, the entanglement entropy between layers is merely the thermal equilibrium entropy for a system with the Hamiltonian $H_{0}$ and inverse temperature $\beta$. Note, however, that for other partitions the entanglement entropy can be nonthermal (see refs. Langlett et al. (2022); Wildeboer et al. (2022) and the example below).
It is worth noting that the reduced state (10) is a steady state of the Lindbladian ${\mathcal{L}}$.
@@
They are referred to as Hamiltonian integrals of motion (HIoM). We assume that a HIoM $Q$ is an eigen operator of the adjoint dissipator and hence of the adjoint Lindladian,
@@
Remarkably, there are broad classes of models where eigen-HIoMs can be explicitly found. The model (11), (12) is one example. There the condition (17) is satisfied for the Hamiltonian as well as for polynomials of arbitrary power constructed from $S^{z}$ sup :
$Q_{1}$ $=H,$ $\Gamma_{1}$ $=\gamma\cosh\frac{\beta}{2},$
$Q_{2}$ $=S^{z}+\frac{N}{2}\tanh\frac{\beta}{2}\cdot{\mathbb{1}},$ $\Gamma_{2}$ $=\gamma\cosh\frac{\beta}{2},$
$Q_{3}$ $=\left(S^{z}\right)^{2}+(N-1)\tanh\frac{\beta}{2}S^{z}$
$+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\cdot{\mathbb{1}},$ $\Gamma_{3}$ $=2\,\gamma\cosh\frac{\beta}{2},$
@@
[…]

### v4

@@ Abstract
We establish a duality between open many-body systems governed by the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) equation and satisfying the detailed balance condition on the one side, and closed bilayer systems with a self-adjoint Hamiltonian on the other side. Under this duality, the identity operator on the open system side maps to a quantum many-body scar of the dual Hamiltonian $\mathcal{H}$. This scar eigenstate has a form of a thermofield double state for a single-body conserved quantity entering the detailed balance conditions. A remarkable feature of this thermofield scar is a tunable single-layer entanglement entropy controlled by the reservoir temperature on the open system side. Further, we identify broad classes of many-body open systems with nontrivial explicit eigen operators $Q$ of the Lindbladian superoperator. The expectation values of the corresponding observables exhibit a simple exponential decay, $\langle Q\rangle_{t}=e^{-\Gamma t}\langle Q\rangle_{0}$, irrespectively of the initial state. Under the above duality, these eigen operators give rise to additional (towers of) scars. Finally, we point out that more general superoperators (not necessarily of the GKSL form) can be mapped to self-adjoint Hamiltonians of bilayer systems harbouring scars, and provide an example thereof.
@@
Remarkably, the scar obtained here turns out to be a thermofield double state with respect to a single-body observable which would be a conserved quantity in the absence of dissipation – thus we adopt the term “thermofield scar”. The thermofield double state first emerged in thermal quantum field theory Umezawa et al. (1982) and further found applications in molecular dynamics Borrelli and Gelin (2021), black hole physics Maldacena (2003); Maldacena and Susskind (2013); Hartman and Maldacena (2013) and quantum simulations Wu and Hsieh (2019); Zhu et al. (2020); Francis et al. (2021). Importantly, the single-layer bipartite entanglement entropy of the thermofield scar is thermal, with the temperature coinciding with the reservoir temperature on the open system side of the duality. In a special case of infinite temperature, the entanglement entropy is maximal and the thermofield scar reduces to the recently discovered Einstein-Podolsky-Rosen (EPR) scar Langlett et al. (2022); Wildeboer et al. (2022). At the same time, for other bipartitions the entanglement entropy of the thermofield scar is zero, which justifies the name “scar” Serbyn et al. (2021); Moudgalya et al. (2022).
@@
The rest of this paper is organized as follows. We start by introducing the duality. Then we show that the duality implies the existence of the thermofield scar. After that we illustrate our general findings in a particular family of spin-$1/2$ models. Further, we demonstrate that in broad classes of dissipative models one can construct explicit eigen operators of the adjoint Lindbladian, and that these operators map into additional scars in dual bilayer closed systems. Then we briefly discuss dualities between bilayer Hamiltonians and superoperators of non-GKSL form. We conclude by discussing our results in relation to other developments and outlining further prospects of our approach.
@@
Here $O_{t}$ is a Heisenberg operator of some observable, $\mathcal{L}$} is a Lindbladian superoperator (Lindbladian), $H$ is the Hamiltonian which is responsible for the coherent part of evolution, and $\cal D$ is the dissipation superoperator (dissipator) to be defined later. The operator $O_{t}$ acts in a Hilbert space $\mathscr{H}$. The expectation value $\langle O\rangle_{t}$ of the observable $O$ evolves in time according to $\langle O\rangle_{t}=\operatorname{\mbox{tr}}\rho_{0}\,O_{t}$, where $\rho_{0}$ is the initial state of the open system.
[…]
@@
First we recall the general form of the GKSL dissipator, ${\mathcal{D}}^{\dagger}O_{t}=\sum_{j}\left(\tilde{L}_{j}^{\dagger}O_{t}\tilde{L}_{j}-(1/2)\{\tilde{L}_{j}^{\dagger}\tilde{L}_{j},O_{t}\}\right)$, where $\{\bullet,\bullet\}$ is an anticommutator and $\tilde{L}_{j}$ are arbitrary operators Gorini et al. (1976); Lindblad (1976); Breuer et al. (2002). The dissipator is said to satisfy the detailed balance condition if it can be written in the form (Alicki and Lendi, 2007, Sec. 1.3.4)
@@
where $\beta$ is a real parameter having the meaning of the inverse reservoir temperature in microscopic derivations of the GKSL equation (Breuer et al., 2002, Sec. 3.3.2), $\gamma>0$ is the dissipative coupling constant and (rescaled) Lindblad operators $L_{j}$ are annihilation-like operators with respect to some $H_{0}$ that commutes with $H$, i.e.
$\beta H_{0},L_{j}]=-\beta L_{j},\quad[H_{0},H]=0.$ (3)
The dissipator (2) can be written in the standard GKSL form by replacing rescaled Lindblad operators $L_{j}$ by conventional Lindblad operators $\tilde{L}_{j}$ according to $\tilde{L}_{2j}=\sqrt{\gamma}e^{\frac{\beta}{4}}L_{j}$ , $\tilde{L}_{2j+1}=\sqrt{\gamma}e^{-\frac{\beta}{4}}L_{j}^{\dagger}$. Note that the detailed balance condition entails the constraint $\tilde{L}_{2j}^{\dagger}=e^{\frac{\beta}{2}}\tilde{L}_{2j+1}$.
The detailed balance condition (2), (3) naturally emerges when deriving the GKSL equation for a system interacting with a thermal reservoir in weak coupling or low density limits (Alicki and Lendi, 2007, Sec. 1.3.4). [^1] In fact, the detailed balance condition has a more general form, which can be obtained by replacing $\beta\rightarrow\beta\omega_{j}$ in the right hand sides of eqs. (2) and (3), where $\omega_{j}$ can be all different (Alicki and Lendi, 2007, Sec. 1.3.4). Further generalizations of the detailed balance in quantum systems are discussed in Fagnola and Umanita (2010). While our construction can be generalized appropriately, we leave such generalizations to further studies. Under this condition, the GKSL equation entails a classical Markov equation with classical detailed balance for populations of the conserved quantity $H_{0}$ (see e.g. (Breuer et al., 2002, Section 3.3.2)).
@@
$+\mathbb{1}\otimes\left(H^{T}-\frac{g}{4}\sum_{j}\left(e^{\frac{\beta}{2}}L_{j}^{T}L_{j}^{*}+e^{-\frac{\beta}{2}}L_{j}^{*}L_{j}^{T}\right)\right)$
$+\frac{g}{2}\sum_{j}\biggl(L_{j}^{\dagger}\otimes L_{j}^{T}+L_{j}\otimes L_{j}^{*}\biggr).$ (6)
@@
It should be stressed that the Wick rotation along, without the similarity transformation (5), does not lead to a self-adjoint dual Hamiltonian unless all Lindblad operators can be chosen to be self-adjoint, as in eq. (4). The latter case was considered in our previous paper Teretenkov and Lychkovskiy (2024). Furthermore, a general Lindbladian can not be mapped to a self-adjoint Hamiltonian by means of any similarity transformation and/or Wick rotation. This can be seen from the fact that such mapping would imply the diagonalizability of the Lindbladian, which is not guaranteed even for a two-level system Lendi (1987). These observations highlight the crucial role of the detailed balance condition for obtaining the duality.
Note that a different choice of the basis in eq. (5) leads to a different duality. In particular, matrix transposition and complex conjugation are basis-dependent operations, therefore $H^{T}$, $L_{j}^{T}$ and $L_{j}^{*}$ entering eq. (6) depend on the basis choice. One consequence of the freedom to chose the basis is discussed in the next section.
The duality (5), (6) is the first main result of the present Letter.
@@
This is the second main result of the Letter.
Figure 2: Entanglement of the thermofield scar (8) with respect to different bipartitions of a bilayer system. (a) For the bipartition that singles out one layer and traces out another one, the entanglement entropy equals the thermal entanglement entropy for the Hamiltonian $H_{0}$ at temperature $\beta$. (b) The entanglement entropy vanishes for a transverse bipartition that does not cut the interlayer bonds (here the underlying basis vectors are assumed to be of product form, see the text for details).
The state (8) is a thermofield double state Umezawa et al. (1982) for $H_{0}$. It possesses a remarkable entanglement structure highly dependent on the bipartition. If one traces out one layer of the thermofield scar (8), see Fig. 2 (a), one obtains a reduced state of the second layer which turns out to be a Gibbs state with respect to $H_{0}$ (Nastase, 2023, Section 32.5), (sup, , Eq. (S8)),
@@
where normalization is omitted. [^2] Curiously, the reduced state (10) is a steady state of the Lindbladian $\mathcal{L}$}. As a consequence, the entanglement entropy between layers is merely the thermal equilibrium entropy for a system with the Hamiltonian $H_{0}$ and inverse temperature $\beta$.
However, for “transverse” bipartitions shown in Fig 2(b), the picture is completely different. The precise value of the entanglement entropy for such bipartitions depends on the choice of the basis underlying the duality. In the extreme case where each basis vector is of product form, the thermofield state (8) factorizes for any transverse bipartition and the entanglement entropy is equal to zero. This is explicitly demonstrated for an exemplary system considered in what follows.
In general, the double layer Hamiltonian (6) is nonintegrable and non-localized (here we assume no disorder on the open system side). Further, zero energy is, in general, in the middle of the spectrum. Therefore the zero energy eigenstate (8) possessing a zero entanglement entropy for certain bipartitions is classified as a scar Serbyn et al. (2021); Moudgalya et al. (2022). We adopt the term “thermofield scar” for this state.
We remark that in some cases, even after the thermofield scar and related towers of special eigenstates discussed below are removed from the spectrum, the rest of the spectrum may not appear completely chaotic for various reasons. In this case the term “scar” applied to the eigenstate (8) may become a misnomer, since scars are supposed to be embedded in the surrounding of chaotic eigenstates. With this caveat in mind, we nevertheless keep using the term “thermofield scar” for the sake of brevity.
A particular case of the thermofield scar (8) with $\beta=0$ has been recently discovered in refs. Langlett et al. (2022); Wildeboer et al. (2022) (without using or recognizing the duality). [^3] See also ref. Cottrell et al. (2019) where a local bilayer Hamiltonian is constructed whose ground state is approximately the thermofield double state. It has been tagged as the EPR scar in ref. Wildeboer et al. (2022), for reasons that will be clarified in the example below. Refs. Langlett et al. (2022); Wildeboer et al. (2022) contain an in-depth analysis of the various features of infinite-temperature thermofield scars, including a numerical demonstration of persistent oscillations of local observables for a quantum quench.
@@
Figure 3: Scars and scarred dynamics for the bilayer system (14) with $N=6$ spins in each layer. The Hamiltonian parameters are $g=1$, $\beta=0.7$, $\omega_{j\,j+1}=1$, $\omega_{j\,j+2}=0.5$ and $\omega_{ij}=0$ otherwise. (a) Rényi entropy $(-\log\operatorname{\mbox{tr}}\rho_{E}^{2})$ of eigenstates with respect to the transverse bipartition (see Fig. 2(b)) in two equal parts ($\rho_{E}$ is the corresponding reduced density matrix of an eigenstate). Highlighted are the thermofield scar $|\mathbb{1}\rangle\rangle_{\beta}$ and the scars $|Q_{n}\rangle\rangle_{\beta}$, $n=1,2,\dots,6$. (b)  Overlaps $|\langle E|\Psi_{0}\rangle|^{2}$ between the eigenstates and the product initial state $\Psi_{0}=|\downarrow\downarrow\dots\downarrow\rangle$ (all spins down in both layers). One can see that this initial state is expanded over the thermofield scar and the scars $|Q_{n}\rangle\rangle_{\beta}$. (c) Persistent oscillations of the squared total spin polarizations $\left(S^{x,y,z}\right)^{2}$ after initialization in the state $\Psi_{0}$.
@@
They are referred to as Hamiltonian integrals of motion (HIoM). We assume that a HIoM $Q$ is an eigen operator of the adjoint dissipator and hence of the adjoint Lindbladian,
@@
Remarkably, there are broad classes of models where eigen-HIoMs can be explicitly found. The model (11), (12) is one example. There the condition (17) is satisfied for a set $Q_{n}$ of eigen-HIoMs that are polynomials constructed from $S^{z}$, $n\leq N$ being the polynomial degree. The first two of them read sup
$Q_{1}$ $=S^{z}+\frac{N}{2}\tanh\frac{\beta}{2}\cdot{\mathbb{1}},$ $\Gamma_{1}$ $=\gamma\cosh\frac{\beta}{2},$
$Q_{2}$ $=\left(S^{z}\right)^{2}+(N-1)\tanh\frac{\beta}{2}S^{z}$
$+\frac{N}{4}\left((N-1)\tanh^{2}\frac{\beta}{2}-1\right)\cdot{\mathbb{1}},$ $\Gamma_{2}$ $=2\,\gamma\cosh\frac{\beta}{2},$
@@
In addition, the single-layer Hamiltonian $H$ is also an eigen-HIoM with $\Gamma=\gamma\cosh\frac{\beta}{2}$.
@@
Explicit expressions for scars of the model (14) dual to eigen-HIoMs (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ) can be found in the Supplement sup . Note that they do not have a strictly vanishing entanglement entropy for the transverse bipartition. In Fig. 3 we show these scars immersed in the spectrum of the model (14). We also demonstrate that, analogously to the infinite-temperature case Langlett et al. (2022), the initial product state $\Psi_{0}=|\downarrow\downarrow\dots\downarrow\rangle$ (all spins down in both layers) has large overlaps with the thermofield scar (8) and the scars from the tower dual to HIoMs (Duality between open systems and closed bilayer systems, and thermofield double states as quantum many-body scars ), which entails persistent oscillations of the total spin projections squared, $\left(S^{\alpha}\right)^{2}(t)$, $\alpha=x,y,z$.
Curiously, the spectrum of the model (14) apparently contains other scars. Two of them, $|P_{1}\rangle=\bigotimes_{j=1}^{N}|\uparrow\,\downarrow\rangle_{j}$ and $|P_{2}\rangle=\bigotimes_{j=1}^{N}|\downarrow,\uparrow\rangle_{j}$ (spins up in one layer, spins down in another layer), are of the product form and do not depend on $\beta$. While they can be understood in terms of the duality sup , in fact they are examples of curious special eigenstates generic for a wider class of bilayer systems beyond the duality sup . We leave deeper exploration of the model (14) for further studies.
@@
Beyond GKSL.  Finally, let us briefly comment on dualities in a more general perspective. The dimension counting tells us that, in fact, an arbitrary self-adjoint bilayer Hamiltonian $\mathcal{H}$ can be mapped to some superoperator $\mathcal{S}$ acting on operators in a Hilbert space corresponding to a single layer. This superoperator is not necessarily of the GKSL form, therefore, at the first sight, such a duality misses much of its physical appeal. [^4] Note that non-Markovian dynamics of open quantum systems can lead to GKSL-like, but non-GKSL generators Hall et al. (2014). However, it can still be very useful, in particular for identifying special eigenstates of $\mathcal{H}$. To illustrate the latter claim, consider a bilayer spin-1/2 Hamiltonian $\mathcal{H}=\sum_{i<j}w_{ij}\left(\,\sigma^{x}_{i}\sigma^{z}_{j}+\tau^{x}_{i}\tau^{z}_{j}\right)+\sum_{j}\sigma^{y}_{j}\tau^{y}_{j}$ (notations are the same as in the Hamiltonian (14)). The infinite-temperature version of the map (5) leads to the superoperator ${\mathcal{S}}O=\{H,O\}-\sum_{j}\sigma^{y}_{j}\,O\,\sigma^{y}_{j}$, where $H=\sum_{i<j}w_{ij}\sigma^{x}_{i}\sigma^{z}_{j}$ and $\{\bullet,\bullet\}$ is the anticommutator. Clearly, $\mathcal{S}$} is not of the GKSL form. However, one easily identifies two eigen operators $Q_{x,z}$ of $\mathcal{S}$}, $Q_{x,z}=\prod_{j}\sigma_{j}^{x,z}$, satisfying ${\mathcal{S}}\,Q_{x,z}=N\,Q_{x,z}$. They correspond to two eigenstates $|Q_{x}\rangle\rangle=\bigotimes_{j}\left(|\uparrow\,\downarrow\rangle_{j}+|\downarrow\,\uparrow\rangle_{j}\right)$ and $|Q_{z}\rangle\rangle=\bigotimes_{j}\left(|\uparrow\,\uparrow\rangle_{j}-|\downarrow\,\downarrow\rangle_{j}\right)$ of the bilayer Hamiltonian $\mathcal{H}$ (notations are the same as in eq. (15)). This example reveals that viewing bilayer systems through the prism of a general duality of type (5) can help identify their features concealed otherwise. Exploring such dualities constitutes a promising direction for further research.
@@
Much of the theory presented here remains valid in the case of time-dependent Hamiltonian and Lindblad operators. In particular, the thermofield state (8) remains a zero energy stationary state of the now-time-dependent dual bilayer Hamiltonian in this case, provided the detailed balance conditions (2) are satisfied at each moment of time. This paves the way to Floquet scars Mizuta et al. (2020); Yarloo et al. (2020); Mukherjee et al. (2020); Moudgalya et al. (2022) in periodically driven bilayer systems, as well as to even more exotic stationary states of bilayer Hamiltonians that depend on time aperiodically. We leave this direction of study for further work.
@@
It can be interesting to explore the freedom to use different bases for defining the duality is yet to be explored. Here we have focused on the simplest case of product state bases; more complex bases will lead to dualities with more complex properties and, in particular, to thermofield scars with more intricate entanglement structure.
Finally, we note that the reported duality might aid quantum simulations of open systems David et al. (2024) and shed new light on topological effects in open systems Niu and Wang (2024).